# Singularity-free cosmology from interactions in the dark sector

### v1

@@ III.2 Parameter Space for Viable Models
In this subsection, we study how the fixed points and their eigenvalues change for $\epsilon=\pm 1$, and for $w_{x}>-1$ and $w_{x}<-1$. Notice that because of the interaction term, and because of the $\rho_{\Lambda}$ term in Eq. (2), the dark energy can be phantom, $x^{\prime}>0$, even for $w_{x}>-1$. We assume $q>0$, meaning that during expansion energy transfers from the dark energy to the dark matter, and we take $10<\mathcal{R}<$ 1.
@@ III.2.1 $w_{x}<-1$, $\epsilon=+1$
For $dS_{2+}$ to have negative eigenvalues and be an attractor when $\epsilon=+1$, we require $-1-\mathcal{R}<w_{x}<-1$ and $q<3/\mathcal{R}$. However, for these parameter ranges, $dS_{1+}$ cannot have positive eigenvalues, and therefore cannot be a repellor. More generally, when $dS_{2+}$ is an attractor, $dS_{1+}$ can only exist at $z<0$. We also note that for these parameter ranges, $dS_{3+}$ cannot be a repellor. Therefore, for $w_{x}<-1$ and $\epsilon=+1$, we cannot have a high energy repellor and a low energy attractor.
@@ III.2.2 $w_{x}<-1$, $\epsilon=-1$
For the eigenvalues of $dS_{2+}$ to be negative when $\epsilon=-1$, we require $w_{x}>-1+\mathcal{R}$. Therefore, $dS_{2+}$ cannot be an attractor for $\epsilon=-1$ and $w_{x}<-1$, and so we do not have have a late-time cosmological constant for these parameter values.
@@ III.2.3 $w_{x}>-1$, $\epsilon=+1$
For the $dS_{2+}$ fixed point to have negative eigenvalues when $\epsilon=+1$, we require $w_{x}>-1$ and $q<3/\mathcal{R}$. However for these parameter values, $dS_{1+}$ cannot have eigenvalues with positive real parts and so cannot be a repellor. More generally, $dS_{1+}$ only exists at $z<0$ for these parameters. The fixed point $dS_{3+}$ can be a repellor for these parameter ranges, however it only exists at $x<0$. Therefore, we cannot achieve our desired topology when $w_{x}>-1$ and $\epsilon=+1$.
@@ III.2.4 $w_{x}>-1$, $\epsilon=-1$
Setting $\epsilon=-1$ and $10<\mathcal{R}<$ 1, we find that $dS_{2+}$ is an attractor when $q<3/\mathcal{R}$ and $w_{x}>-1+|\epsilon|\mathcal{R}$. These parameter constraints are sufficient for $dS_{1+}$ to exist at $z>0$ and be a repellor. Depending on the specific value of $q$, the stability of $dS_{1+}$ can change. It is possible for the eigenvalues to be real and positive such that the fixed point is a repelling node, and it is also possible for the eigenvalues to be complex with positive real parts such that the fixed point is a spiral repellor. There also exits a limiting value between these two cases, such that the eigenvalues are real, positive and the same, meaning the repellor is an improper node. With these combinations of parameters, we find that $dS_{3+}$ exists at $x>0$ and is a saddle fixed point, which always has a larger $x$ value than $dS_{2+}$, and always has a smaller $x$-value than $dS_{1+}$. For these parameters, the dark energy violates the NEC when $x<\mathcal{R}$ and when $x>1+w_{x}$, and so has phantom behavior. Therefore, a high energy repellor and an attractor at low energy exist when $\epsilon=-1$ and $w_{x}>-1$, however we need to understand whether all trajectories emerge from $dS_{1+}$ during expansion and evolve towards $dS_{2+}$. From now on, we fix $\epsilon=-1$, $-1+\mathcal{R}<w_{x}<0$ and $0<q<3/\mathcal{R}$.
@@ III.3 Behavior at infinity
Thus far, we have discussed the fixed points in the phase plane for $x$ and $z$, however to understand the global behavior we introduce compactified variables. In the system above, $y$ takes values in the range $-\infty<y<\infty$, and we are interested in values of $x$ and $z$ in the range $0<x,z<\infty$, such that the energy densities of the dark components are always positive. In order to understand the dynamics at infinity, we compactify our variables as follows,
@@
where $X=Z=1$ corresponds to $x,z\rightarrow\infty$, and $Y=\pm 1$ corresponds to $y\rightarrow\pm\infty$. For $z$, it is obvious that $z$ is always greater than zero from Eq. (10). Along $x=0$, $x^{\prime}>0$, meaning for any positive initial $x$ value, $x$ will remain positive if we evolve the system forwards in time. However, $x$ can be negative and cross $x=0$ from below, therefore we need to set variable ranges such that we only consider trajectories where $x$ is always positive. We therefore analyze our system between $10<X,Z<$ 1, and $-1<Y<1$. The system of ODEs in Eq. (9) - (11) in compact variables become,
@@
We can also write Eq.s (15) and (16) in compact form as
@@
which have critical values/ singularities at $X=1$ and $Z=1$, corresponding to $x\to\infty$ and $z\to\infty$ respectively. These critical points are shown in Table 3, and a schematic diagram of the fixed points, critical points and separatrices in the 2-D system is shown in Fig. 1. It is especially important to understand the dynamics near $(Z,X)=(1,1)$, and in particular we want to know whether a separatrix exists between $dS_{1+}$ and $S_{2+}$. If such a separatrix does exist, then there will be trajectories trapped between it and the $X=1$ line that evolve to $S_{2+}$. On the other hand, if such a separatrix does not exist, then $S_{2+}$ will be a generalized saddle point and all trajectories will avoid the singularity and evolve to $dS_{2+}$.
@@
$S_{1\pm}$ 00 $1$
$S_{2\pm}$ $1$ $1$
$S_{3\pm}$ $1$ 00
Table 3: The critical points that arise due to making the system compact. $S_{\pm}$ denotes a singularity with infinite expansion (+) or contraction (-). Note that in the $Z$-$X$ phase spaces, only the $S_{+}$ critical points are plotted as we have fixed $Y>0$ so that the trajectories are expanding.
First, we multiply Eq.s (23) and (24) by $(1-X)(1-Z)$ which we can view as a rescaling of the vector field or a reparameterization, valid in the unit square $10<X,Z<$ 1. The singular point $(1,1)$ becomes a fixed point, and we can then Taylor expand around $(1,1)$; the linear part vanishes, and the equations to second order become
@@
through the point $(1,1)$ which naturally leads to two cases: when $q<3$, this invariant line lies in the unit square $10<X,Z<$ 1. The flow restricted to this line has the point $(1,1)$ as an attractor, and so there is a sector between this line and the $X=1$ line where all trajectories go onto $(Z,X)=(1,1)$ in the future. This invariant line, followed backwards in time, goes to $dS_{1+}$. Therefore, when $q<3$ a separatrix between $dS_{1+}$ and $S_{2+}$ exists (see Figure 2(a) for a representative example).
On the other hand, when $q>3$ the invariant line does not enter the unit square and it appears all trajectories avoid the singularity at $(1,1)$, ultimately evolving towards $dS_{2+}$ (see Figure 7 for a representative example). We confirm this by writing $X=X(Z)$ and combining the equations in (25) to give a homogeneous ODE for $dX/dZ$ with solution
@@
where $C>0$ is derived from the initial conditions and $a=\frac{3}{q}-1<0$ for $q>3$. As $Z$ tends to 1 from below, the solution grows towards $X=1$, however it reaches a maximum at $Z=1-\left(\tfrac{a+1}{C}\right)^{1/a}$ and then starts to decrease; as such no trajectories in the unit square go onto $(Z,X)=(1,1)$ when $q>3$. Alternatively we could consider (15) and (16) in the ‘high energy regime’ where we keep only the quadratic terms:
@@
Figure 1: A schematic diagram of the fixed points and critical points in the 2-D $Z$-$X$ phase space. The solid black line shows the separatrix between $dS_{1+}$ and $dS_{3+}$ which exists for all values of $q$. The dashed black line shows the separatrix between $dS_{1+}$ and $S_{2+}$ which only exists for $q<3$ (see Sec. III.3 for an explanation).
@@ IV Dynamics with $q<3$
In this section, we analyze the dynamics when $q<3$. For $q<3$, there are models that become singular, however there are trajectories that are nonsingular and qualitatively interesting. We fix $\mathcal{R}=0.01$, however for realistic models we would expect $10^{-120}<\mathcal{R}<10^{-60}$. Ultimately, the dynamics is not affected by how small $\mathcal{R}$ is; as $\mathcal{R}\rightarrow 0$, $dS_{2+}$ just moves closer to the origin. In reality we could fix $w_{x}$ such that the system depends on one parameter and only change $q$ to show the full range of dynamics. However, we change both $w_{x}$ and $q$ to ensure the phase spaces are readable, while maintaining the same topology.
It is clear from Table 2 that there are three possible types of stability character for the repellor $dS_{1+}$: a repelling node, an improper node or a spiral repellor. In all three cases the dynamics is the same. In the following, we show the phase spaces where $dS_{1+}$ is a spiral repellor, however we present examples of the improper node and repelling node cases in the Appendix.
@@
$S_{2+}$ Non-simple ($q<3$) or generalized saddle ($q>3$) Black
@@
Table 4: The stability and color of the fixed points and critical points of the $Z$-$X$ system in the expanding case ($Y>0$), with $\epsilon=-1$, $-1+\mathcal{R}<w_{x}<0$ and $10<\mathcal{R}<$ 1. $dS_{+}$ denotes an expanding (+) de-Sitter universe ($x^{\prime}=z^{\prime}=0$), and $S_{+}$ denotes a singularity with infinite expansion (+).
@@
Table 5: The colors of different features in the 2-D $Z$-$X$ phase space.
In the $Z$-$X$ phase spaces two separatrices exist (see Fig. 1). The first is the repellor-saddle separatrix, which joins the $dS_{1+}$ repellor and the $dS_{3+}$ saddle, and the second is the repellor-singularity separatrix, which joins $dS_{1+}$ and the singularity $S_{2+}$. These separatrices separate two types of trajectories. The first evolve to the right of these separatrices, which are non-singular models. They emerge during expansion from the non-singular de Sitter fixed point $dS_{1+}$ and expand toward the late-time attractor $dS_{2+}$, where the dark energy asymptotically approaches a cosmological constant and the dark matter energy density tends to zero. Some trajectories have decreasing dark energy density ($\dot{X}<0$) as they approach $dS_{2+}$, and some have phantom behavior with increasing dark energy density ($\dot{X}>0$), which can be seen in Fig 2(b). Trajectories to the left of the separatrices also expand from $dS_{1+}$, but evolve towards $S_{2+}$.
There are three sub-cases for the $Z$-$X$ plane which we present. The topology of the phase space is the same, but they represent physically different models. The difference in each case is whether the repellor-saddle separatrix and the zero acceleration $a^{\prime\prime}=0$ curve intersect twice, touch or do not intersect. The zero acceleration curve is found by setting Eq. (12) equal to zero. In compact variables, this equation is
@@
which we plot in red in the 2-D phase spaces. In all cases, either the $a^{\prime\prime}=0$ curve crosses the $Z=0$ axes below the $dS_{2+}$ attractor, or it does not cross the $X$-axis at all, therefore all trajectories have late-time acceleration. In the full 3-D system, where we include the equation for $Y$, Eq. (30) represents a curve of static Einstein models along $Y=0$. In Sec. IV.2, we only plot first integral surfaces in the $X$-$Y$-$Z$ phase spaces, set by a specific $Z_{0}$ and $X_{0}$, which correspond to a single trajectory in the 2-D phase space. If the trajectory touches or intersects the $a^{\prime\prime}=0$ curve in the $Z$-$X$ phase space, then the first integral surface touches or intersects the zero acceleration surface in the corresponding 3-D phase space. Einstein fixed points exist in the 3-D phase spaces where these two surfaces touch or intersect at $Y=0$. The physically interesting models are the ones which have a decelerated period, which are represented by the trajectories which intersect the $a^{\prime\prime}=0$ curve twice in the 2-D phase space, and correspond to a first integral surface in 3-D which have two Einstein fixed points along $Y=0$.
In the following, we first present the three cases for the 2-D phase spaces. The stability of the fixed points and critical points in the expanding ($Y>0$) $Z$-$X$ phase spaces, along with their color in the plots, are given in Table 4, and the color scheme for the different types of curves are given in Table 5.

### v2

@@ III.2 Parameter space for viable models
In this subsection, we study how the fixed points and their eigenvalues change for $\epsilon=\pm 1$, and for $w_{x}>-1$ and $w_{x}<-1$. Notice that because of the interaction term, and because of the $\rho_{\Lambda}$ term in Eq. (2), the dark energy can be phantom, $x^{\prime}>0$, even for $w_{x}>-1$. We assume $q>0$, meaning that during expansion energy transfers from the dark energy to the dark matter, and we take $0<\mathcal{R}<1$.
@@ III.2.1 $w_{x}<-1$, $\epsilon=+1$
For $dS_{2+}$ to have negative eigenvalues and be an attractor when $\epsilon=+1$, we require $1-\mathcal{R}<w_{x}<-1$ and $q<3/\mathcal{R}$. However, for these parameter ranges, $dS_{1+}$ cannot have positive eigenvalues, and therefore cannot be a repellor. More generally, when $dS_{2+}$ is an attractor, $dS_{1+}$ can only exist at $z<0$. We also note that for these parameter ranges, $dS_{3+}$ cannot be a repellor. Therefore, for $w_{x}<-1$ and $\epsilon=+1$, we cannot have a high energy repellor and a low energy attractor.
@@ III.2.2 $w_{x}<-1$, $\epsilon=-1$
For the eigenvalues of $dS_{2+}$ to be negative when $\epsilon=-1$, we require $w_{x}>-1+\mathcal{R}$. Therefore, $dS_{2+}$ cannot be an attractor for $\epsilon=-1$ and $w_{x}<-1$, and so we do not have a late-time cosmological constant for these parameter values.
@@ III.2.3 $w_{x}>-1$, $\epsilon=+1$
For the $dS_{2+}$ fixed point to have negative eigenvalues when $\epsilon=+1$, we require $w_{x}>-1$ and $q<3/\mathcal{R}$. However for these parameter values, $dS_{1+}$ cannot have eigenvalues with positive real parts and so cannot be a repellor. More generally, $dS_{1+}$ only exists at $z<0$ for these parameters. The fixed point $dS_{3+}$ can be a repellor for these parameter ranges; however, it only exists at $x<0$. Therefore, we cannot achieve our desired topology when $w_{x}>-1$ and $\epsilon=+1$.
@@ III.2.4 $w_{x}>-1$, $\epsilon=-1$
Setting $\epsilon=-1$ and $0<\mathcal{R}<1$, we find that $dS_{2+}$ is an attractor when $q<3/\mathcal{R}$ and $w_{x}>-1+|\epsilon|\mathcal{R}$. These parameter constraints are sufficient for $dS_{1+}$ to exist at $z>0$ and be a repellor. Depending on the specific value of $q$, the stability of $dS_{1+}$ can change. It is possible for the eigenvalues to be real and positive such that the fixed point is a repelling node, and it is also possible for the eigenvalues to be complex with positive real parts such that the fixed point is a spiral repellor. There also exits a limiting value between these two cases, such that the eigenvalues are real, positive and the same, meaning the repellor is an improper node. With these combinations of parameters, we find that $dS_{3+}$ exists at $x>0$ and is a saddle fixed point, which always has a larger $x$ value than $dS_{2+}$, and always has a smaller $x$-value than $dS_{1+}$. For these parameters, the dark energy violates the NEC when $x<\mathcal{R}$ and when $x>1+w_{x}$, and so has phantom behavior. Therefore, a high energy repellor and an attractor at low energy exist when $\epsilon=-1$ and $w_{x}>-1$; however, we need to understand whether all trajectories emerge from $dS_{1+}$ during expansion and evolve towards $dS_{2+}$. From now on, we fix $\epsilon=-1$, $1+\mathcal{R}<w_{x}<0$ and $0<q<3/\mathcal{R}$.
@@ III.3 Behavior at infinity
Thus far, we have discussed the fixed points in the phase plane for $x$ and $z$; however, to understand the global behavior we introduce compactified variables. In the system above, $y$ takes values in the range $\infty<y<\infty$, and we are interested in values of $x$ and $z$ in the range $0<x,z<\infty$, such that the energy densities of the dark components are always positive. We note that in Paper I Burkmar and Bruni (2023), we considered the dynamics in the range $\mathcal{R}<x<1$; if we had considered $x<\mathcal{R}$ or $x>1$, trajectories in these regions would have always been phantom, and therefore qualitatively unrealistic. At high energy, the interaction term decreases $x$, and is able to counter the phantom runaway behavior caused by the quadratic term in Eq. (9), and at low energy, trajectories which were nonphantom in the past can be phantom at late times, unlike in Paper I Burkmar and Bruni (2023). Therefore, in this paper we study $x$ in the range $0<x<\infty$.
In order to understand the dynamics at infinity, we compactify our variables as follows:
@@
where $X=Z=1$ corresponds to $x,z\rightarrow\infty$, and $Y=\pm 1$ corresponds to $y\rightarrow\pm\infty$. For $z$, it is obvious that $z$ is always greater than zero from Eq. (11). Along $x=0$, $x^{\prime}>0$, meaning for any positive initial $x$ value, $x$ will remain positive if we evolve the system forward in time. However, $x$ can be negative and cross $x=0$ from below, therefore we need to set variable ranges such that we only consider trajectories where $x$ is always positive. We therefore analyze our system between $0<X,Z<1$, and 1<Y<1-1<Y<1. The system of ODEs in Eqs. (10) – (12) in compact variables becomes
@@
We can also write Eqs. (16) and (17) in compact form as
@@
which have critical values/ singularities at $X=1$ and $Z=1$, corresponding to $x\to\infty$ and $z\to\infty$ respectively. These critical points are shown in Table 3, and a schematic diagram of the fixed points, critical points and separatrices in the two-dimensional system is shown in Fig. 1. It is especially important to understand the dynamics near $(Z,X)=(1,1)$, and in particular we want to know whether a separatrix exists between $dS_{1+}$ and $S_{2+}$. If such a separatrix does exist, then there will be trajectories trapped between it and the $X=1$ line that evolve to $S_{2+}$. On the other hand, if such a separatrix does not exist, then $S_{2+}$ will be a generalized saddle point and all trajectories will avoid the singularity and evolve to $dS_{2+}$.
@@
$S_{1\pm}$ 0 11
$S_{2\pm}$ 11 11
$S_{3\pm}$ 11 0
Table 3: Critical points that arise due to making the system compact. $S_{\pm}$ denotes a singularity with infinite expansion (+) or contraction (-). Note that in the $Z$-$X$ phase spaces, only the $S_{+}$ critical points are plotted as we have fixed $Y>0$ so that the trajectories are expanding.
First, we multiply Eqs. (24) and (25) by (1X)(1Z)(1-X)(1-Z) which we can view as a rescaling of the vector field or a reparametrization, valid in the unit square $0<X,Z<1$. The singular point $(1,1)$ becomes a fixed point, and we can then Taylor expand around $(1,1)$; the linear part vanishes, and the equations to second order become
@@
through the point $(1,1)$ which naturally leads to two cases: when $q<3$, this invariant line lies in the unit square $0<X,Z<1$. The flow restricted to this line has the point $(1,1)$ as an attractor, and so there is a sector between this line and the $X=1$ line where all trajectories go onto $(Z,X)=(1,1)$ in the future. This invariant line, followed backward in time, goes to $dS_{1+}$. Therefore, when $q<3$ a separatrix between $dS_{1+}$ and $S_{2+}$ exists [see Fig. 2(2(a)) for a representative example].
On the other hand, when $q>3$ the invariant line does not enter the unit square and it appears all trajectories avoid the singularity at $(1,1)$, ultimately evolving towards $dS_{2+}$ (see Fig. 7 for a representative example). We confirm this by writing $X=X(Z)$ and combining the equations in (26) to give a homogeneous ODE for $dX/dZ$ with solution
@@
where $C>0$ is derived from the initial conditions and $a=\frac{3}{q}-1<0$ for $q>3$. As $Z$ tends to 1 from below, the solution grows towards $X=1$; however, it reaches a maximum at $Z=1-\left(\tfrac{a+1}{C}\right)^{1/a}$ and then starts to decrease; as such no trajectories in the unit square go onto $(Z,X)=(1,1)$ when $q>3$. Alternatively we could consider Eqs. (16) and (17) in the “high energy regime” where we keep only the quadratic terms
@@
Figure 1: Schematic diagram of the fixed points and critical points in the two-dimensional $Z$-$X$ phase space. The solid black line shows the separatrix between $dS_{1+}$ and $dS_{3+}$ which exists for all values of $q$. The dashed black line shows the separatrix between $dS_{1+}$ and $S_{2+}$ which only exists for $q<3$ (see Sec. III.3 for an explanation).
@@ IV Dynamics with $q<3$
In this section, we analyze the dynamics when $q<3$. For $q<3$, there are models that become singular; however, there are trajectories that are nonsingular and qualitatively interesting. We fix $\mathcal{R}=0.01$; however, for realistic models we would expect $\mathcal{R}$ to be much smaller if $\rho_{*}$ is between Planck and inflationary energy scales, and $\rho_{\Lambda}$ is close to the observed value of $\Lambda$. In Sec. II, we found a realistic range of $\mathcal{R}$ to be $10^{-120}<\mathcal{R}<10^{-111}$. Ultimately, the dynamics are not affected by how small $\mathcal{R}$ is; as $\mathcal{R}\rightarrow 0$, $dS_{2+}$ just moves closer to the origin. In reality we could fix $w_{x}$ such that the system depends on one parameter and only change $q$ to show the full range of dynamics. However, we change both $w_{x}$ and $q$ to ensure the phase spaces are readable, while maintaining the same topology.
It is clear from Table 2 that there are three possible types of stability character for the repellor $dS_{1+}$: a repelling node, an improper node or a spiral repellor. In all three cases the dynamics is the same. In the following, we show the phase spaces where $dS_{1+}$ is a spiral repellor; however, we present examples of the improper node and repelling node cases in the Appendix.
@@
$S_{2+}$ Nonsimple ($q<3$) or generalized saddle ($q>3$) Black
@@
Table 4: Stability and color of the fixed points and critical points of the $Z$-$X$ system in the expanding case ($Y>0$), with $\epsilon=-1$, $1+\mathcal{R}<w_{x}<0$ and $0<\mathcal{R}<1$. $dS_{+}$ denotes an expanding (+) de Sitter universe ($x^{\prime}=z^{\prime}=0$), and $S_{+}$ denotes a singularity with infinite expansion (+).
@@
Table 5: Colors of different features in the two-dimensional $Z$-$X$ phase space.
In the $Z$-$X$ phase spaces two separatrices exist (see Fig. 1). The first is the repellor-saddle separatrix, which joins the $dS_{1+}$ repellor and the $dS_{3+}$ saddle, and the second is the repellor-singularity separatrix, which joins $dS_{1+}$ and the singularity $S_{2+}$. These separatrices separate two types of trajectories. The first evolve to the right of these separatrices, which are nonsingular models. They emerge during expansion from the nonsingular de Sitter fixed point $dS_{1+}$ and expand toward the late-time attractor $dS_{2+}$, where the dark energy asymptotically approaches a cosmological constant and the dark matter energy density tends to zero. Some trajectories have decreasing dark energy density ($\dot{X}<0$) as they approach $dS_{2+}$, and some have phantom behavior with increasing dark energy density ($\dot{X}>0$), which can be seen in Fig. 2(2(b)). Trajectories to the left of the separatrices also expand from $dS_{1+}$, but evolve towards $S_{2+}$.
There are three subcases for the $Z$-$X$ plane which we present. The topology of the phase space is the same, but they represent physically different models. The difference in each case is whether the repellor-saddle separatrix and the zero acceleration $a^{\prime\prime}=0$ curve intersect twice, touch or do not intersect. The zero acceleration curve is found by setting Eq. (13) equal to zero. In compact variables, this equation is
@@
which we plot in red in the two-dimensional phase spaces. In all cases, either the $a^{\prime\prime}=0$ curve crosses the $Z=0$ axes below the $dS_{2+}$ attractor, or it does not cross the $X$-axis at all, therefore all trajectories have late-time acceleration. In the full three-dimensional system, where we include the equation for $Y$, Eq. (31) represents a curve of static Einstein models along $Y=0$. In Sec. IV.2, we only plot first integral surfaces in the $X$-$Y$-$Z$ phase spaces, set by a specific $Z_{0}$ and $X_{0}$, which correspond to a single trajectory in the two-dimensional phase space. If the trajectory touches or intersects the $a^{\prime\prime}=0$ curve in the $Z$-$X$ phase space, then the first integral surface touches or intersects the zero acceleration surface in the corresponding three-dimensional phase space. Einstein fixed points exist in the three-dimensional phase spaces where these two surfaces touch or intersect at $Y=0$. The physically interesting models are the ones which have a decelerated period, which are represented by the trajectories which intersect the $a^{\prime\prime}=0$ curve twice in the two-dimensional phase space, and correspond to a first integral surface in three dimensions which have two Einstein fixed points along $Y=0$.
In the following, we first present the three cases for the two-dimensional phase spaces. The stability of the fixed points and critical points in the expanding ($Y>0$) $Z$-$X$ phase spaces, along with their color in the plots, are given in Table 4, and the color scheme for the different types of curves are given in Table 5.